# Effect-Transparent Governance for AI Workflow Architectures: Semantic Preservation, Expressive Minimality, and Decidability Boundaries

### 01 May 2026 v1

@@ 1 Introduction
This paper formalizes and mechanically verifies an architectural alternative: effect-level structural governance, extending our prior mechanized development McCann (2026d) with semantic transparency, expressive minimality, and decidability boundary results. Instead of modifying internal computation or post-processing outputs, governance is imposed at the effect boundary of the system. Programs are modeled as interaction trees over a directive event type. Governance is a wrapper over event interpretation in the tradition of algebraic effect handlers Plotkin and Pretnar (2009): it mediates each directive with structural checks (e.g. trust and capability predicates) and governance-only instrumentation (e.g. provenance and observability).
@@
This paper focuses on expressiveness and semantic transparency results. It builds on the mechanized safety and invariance proofs in McCann (2026d) and extends the structural governance criterion introduced in McCann (2026e). The algebraic semantics paper McCann (2026a) lifts these system-specific results to a parametric framework, showing that any system satisfying three axioms inherits the safety and transparency properties proved here. The certified purity paper McCann (2026b) addresses the practical enforcement of the pure module constraint that the safety theorems assume. The distribution provenance paper McCann (2026c) extends governance to the supply chain with cryptographic provenance.
@@ 3 Background
We recall the interaction-tree model and the governance wrapper construction, introduced in McCann (2026d). All definitions are mechanized in Rocq using Interaction Trees Xia et al. (2020); Zakowski et al. (2021).
@@ 10 Discussion and Trust Boundary
Our theorems establish properties of a formal model: interaction-tree programs, a governance wrapper, and a safety predicate. The guarantees hold relative to a standard trusted computing base: the Rocq kernel (type-checking the proofs), the Interaction Trees library (providing the coinductive framework), and any runtime that faithfully implements the $\mathcal{G}$ operator and directive interpretation. This TCB is comparable to that of verified compilers Leroy (2009) and verified OS kernels Klein et al. (2009): formal properties hold of the model, and the deployment gap is bridged by testing. In our case, the connection from model to a deployed runtime is validated via a verified interpreter specification tested against over 70,000 randomly generated directive sequences with zero disagreements (included in the artifact). Effect separation for the executor layer is developed in McCann (2026b).
@@ 11 Related Work
Prior frameworks propose properties to verify Seshia et al. (2016); Dalrymple and others (2024); we deliver a mechanized development with semantic transparency and expressive minimality.
@@
The governance model, safety predicate, and Turing completeness proof are given in McCann (2026d). The structural argument for why behavioral governance fails at the effect level is developed in McCann (2026e). Effect separation for the executor layer is formalized in McCann (2026b). The present paper extends these with semantic transparency, expressive minimality, and a decidability boundary with compositional closure.

### 05 May 2026 v2

@@ 1 Introduction
This paper formalizes and mechanically verifies an architectural alternative: effect-level structural governance, extending our prior mechanized development McCann (2026b) with semantic transparency, expressive minimality, and decidability boundary results. Instead of modifying internal computation or post-processing outputs, governance is imposed at the effect boundary of the system. Programs are modeled as interaction trees over a directive event type. Governance is a wrapper over event interpretation in the tradition of algebraic effect handlers Plotkin and Pretnar (2009): it mediates each directive with structural checks (e.g. trust and capability predicates) and governance-only instrumentation (e.g. provenance and observability).
@@
This paper focuses on expressiveness and semantic transparency results. It builds on the mechanized safety and invariance proofs in McCann (2026b) and extends the structural governance criterion introduced in McCann (2026e). The algebraic semantics paper McCann (2026a) lifts these system-specific results to a parametric framework, showing that any system satisfying three axioms inherits the safety and transparency properties proved here. The certified purity paper McCann (2026d) addresses the practical enforcement of the pure module constraint that the safety theorems assume. The distribution provenance paper McCann (2026c) extends governance to the supply chain with cryptographic provenance.
@@ 3 Background
We recall the interaction-tree model and the governance wrapper construction, introduced in McCann (2026b). All definitions are mechanized in Rocq using Interaction Trees Xia et al. (2020); Zakowski et al. (2021).
@@ 10 Discussion and Trust Boundary
Our theorems establish properties of a formal model: interaction-tree programs, a governance wrapper, and a safety predicate. The guarantees hold relative to a standard trusted computing base: the Rocq kernel (type-checking the proofs), the Interaction Trees library (providing the coinductive framework), and any runtime that faithfully implements the $\mathcal{G}$ operator and directive interpretation. This TCB is comparable to that of verified compilers Leroy (2009) and verified OS kernels Klein et al. (2009): formal properties hold of the model, and the deployment gap is bridged by testing. In our case, the connection from model to a deployed runtime is validated via a verified interpreter specification tested against over 70,000 randomly generated directive sequences with zero disagreements (included in the artifact). Effect separation for the executor layer is developed in McCann (2026d).
@@ 11 Related Work
Prior frameworks propose properties to verify Seshia et al. (2016); Dalrymple et al. (2024); we deliver a mechanized development with semantic transparency and expressive minimality.
@@
The governance model, safety predicate, and Turing completeness proof are given in McCann (2026b). The structural argument for why behavioral governance fails at the effect level is developed in McCann (2026e). Effect separation for the executor layer is formalized in McCann (2026d). The present paper extends these with semantic transparency, expressive minimality, and a decidability boundary with compositional closure.